# Ola: Pushing the Frontiers of Omni-Modal Language Model with Progressive Modality Alignment

### v2

@@ A.3 Data Collection Details
We provide details on collecting video-audio relevant data. Our data comes from two sources with high-quality raw videos: LLaVA-Video-178k [83], which contains 178k raw videos, and FineVideo [18], which contains 42k raw videos. For the open-sourced video data from LLaVA-Video-178k, we first use the Whisper [58] model to generate subtitles. We find that the videos include content in other languages and videos without valid audio, so we design a filtering method for better results. Specifically, we first assess the ratio of English words in the generated subtitles and discard those with a lower ratio, indicating subtitles in other languages. We also discard extremely short subtitles. Then, we use a large language model, Qwen-2.5-72B, to further filter the subtitles. The model is asked to identify meaningless sentences with the following prompt: ”I will give you a subtitle generated from a video. Identify whether the subtitle is complete, fluent, and informative. Answer directly with yes or no and do not add other explanations.” After this procedure, we gathered 41k valid videos. For the videos in FineVideo, as they are already well-processed, we directly use the subtitles for the following steps. We utilize Qwen-2-72B to generate audio-relevant question-answer pairs based on the given videos and subtitles. The prompt for the vision-language model is: ”Please generate at least three questions and answers based on the information in the subtitle. You can refer to the video for additional context. The questions and answers must be highly relevant to the subtitle and video and should not include fabricated content.” We then generate 243k cross-modal video-audio data points from the 81k collected videos. This data is used for stage 3 training for omni-modal alignment.
Appendix B More Showcases
B.1 Text and Audio Understanding

### v3

@@ A.3 Data Collection Details
In the ViT pre-training phase, we employ Qwen2.5-0.5B as the language interface model to establish cross-modal alignment between visual and textual representations. The alignment is optimized via cross-entropy loss, which measures the discrepancy between the predicted and ground-truth token distributions over the joint embedding space. To curate the training data, we aggregate approximately 10M multimodal samples from three primary domains: (1) OCR-intensive datasets (e.g., scene text recognition), (2) visual grounding corpora with region-text pairs, and (3) general vision-language tasks (e.g., image captioning and VQA). For parameter-efficient adaptation, we integrate a Low-Rank Adaptation (LoRA) module into the LLM component, updating only the low-rank decomposition matrices while freezing the original LLM weights. In contrast, all parameters of the ViT encoder are fine-tuned end-to-end to preserve the rich visual representations learned during pre-training.
For the instruction pre-training phase, we curate a multimodal dataset of 20 million image-text pairs from a combination of open-sourced and proprietary in-house datasets, including images from LAION-5B [60], COYO-700M, Conceptual Captions v3, etc, and knowledge-intensive sources from Wikipedia. To further construct the instruction-quality level data, we use GPT-4o and Gemini-Pro to optimize the question and answer pairs to conduct complexity escalation with multi-hop reasoning requirements and quality unification through style transfer to maintain consistency. We scale up the training batch size to 2048 during the pre-training stage to smooth the gradient and continuously train the model.
Furthermore, we provide details on collecting video-audio relevant data. Our data comes from two sources with high-quality raw videos: LLaVA-Video-178k [79], which contains 178k raw videos, and FineVideo [17], which contains 42k raw videos. For the open-sourced video data from LLaVA-Video-178k, we first use the Whisper [55] model to generate subtitles. We find that the videos include content in other languages and videos without valid audio, so we design a filtering method for better results. Specifically, we first assess the ratio of English words in the generated subtitles and discard those with a lower ratio, indicating subtitles in other languages. We also discard extremely short subtitles. Then, we use a large language model, Qwen-2.5-72B, to further filter the subtitles. The model is asked to identify meaningless sentences with the following prompt: ”I will give you a subtitle generated from a video. Identify whether the subtitle is complete, fluent, and informative. Answer directly with yes or no and do not add other explanations.” After this procedure, we gathered 41k valid videos. For the videos in FineVideo, as they are already well-processed, we directly use the subtitles for the following steps. We utilize Qwen-2-72B to generate audio-relevant question-answer pairs based on the given videos and subtitles. The prompt for the vision-language model is: ”Please generate at least three questions and answers based on the information in the subtitle. You can refer to the video for additional context. The questions and answers must be highly relevant to the subtitle and video and should not include fabricated content.” We then generate 243k cross-modal video-audio data points from the 81k collected videos. This data is used for stage 3 training for omni-modal alignment.
Appendix B More Analysis
[…]
Appendix C More Showcases
C.1 Text and Audio Understanding
@@ C.2 Video Understanding
Appendix D Limitations
In this section, we discuss the limitations of our work. Ola focuses solely on text generation for omni-model understanding and does not explore audio generation. This approach limits the model’s ability to fully leverage multimodal inputs, potentially restricting its applicability in scenarios where audio generation is crucial. Future work could address this by incorporating audio generation capabilities to enhance the model’s comprehensive understanding across different modalities.